Two persons A and B start at the same time from two palces 20km apart. If they walk in the same direction , A over takes B in 10 hours and if they walk in opposite direction , they meet in two hours . Find the rates at which the walk
Answer by venugopalramana(3286)

LET THEIR SPEEDS BE A AND B .
WHEN THEY WALK IN SAME DIRECTION THEIR RELATIVE SPEED IS A-B.
DISTANCE =20 KM.
TIME NEEDED TO MEET=20/(A-B) = 10..OR A-B=20/10=2.........................I
WHEN THEY WALK IN OPPOSITE DIRECTION THEIR RELATIVE SPEED IS A+B
DISTANCE =20 KM.
TIME NEEDED TO MEET=20/(A+B)=2...OR...A+B=20/2=10........................II
EQN.I+EQN.II GIVES
2A=12...OR...A=6
SO B=10-6=4......